THE RADIAL HAMILTONIANS FOR THE A^{1}Σ^{+} and X^{1}Σ^{+} STATES OF LiH and Li_{2}

\underline{JOHN A. COXON}, CAMERON S. DICKINSON and TODD C. MELVILLE, Department of Chemistry, Dalhousie University, Halifax, NS, Canada B3H 4J3.

Extensive collections of literature data involving the A^{1}Σ^{+} and X^{1}Σ^{+} states of LiH and the A^{1}Σ_{u}^{+} and X^{1}Σ_{g}^{+} states of Li_{2} have been employed in direct least-squares fits of the radially dependent Hamiltonians. Vibration-rotational and pure rotational transition energies for the ground X^{1}Σ^{+} state of ^{6}LiH, ^{7}LiH, ^{6}LiD and ^{7}LiD, along with the entire collection of A^{1}Σ^{+} − X^{1}Σ^{+} emission and absorption data for the four isotopomers, are used in the fit for LiH; the data set for Li_{2} consists of more than 6000 A^{1}Σ_{u}^{+} − X^{1}Σ_{g}^{+} absorption lines. The Born-Oppenheimer potentials are represented by a modified Lennard-Jones function that provides for correct behaviour in the near dissociation, long-range region of the potential. The LiH data set extends over ≈ 65% and 80% of the well depths of the A^{1}Σ^{+} and X^{1}Σ^{+} states, respectively, while the Li_{2} data covers ≈ 70% and 99.97% of the well depths (extending out to more than 14 Å in the ground X^{1}Σ_{g}^{+} state). The fits take full account of Born- Oppenheimer breakdown effects, and reproduce the measured line positions to within their estimated accuracies.